# Quantum geometry and local moment swapover in correlated graphene heterostructures

### v1

@@ Abstract
Figure 4: Decorated graphene with smaller impurity density. (a) Unit cell of decorated graphene with 1/8 impurity X coverage as in Fig. 3a. The index of sublattice sites Ai, Bi refer to distance from the impurtiy X. (b) Site-resolved weight $w_{m}=\sum_{\bm{k}}|w_{\bm{k}m}|^{2}/N_{\bm{k}}$ ($m\in\{\mathrm{A}_{i},\mathrm{B}_{i},\mathrm{X}\}$) and minimal Wannier spread $\Omega_{\mathrm{I}}=\operatorname{\mathrm{Tr}}g_{ii}(\bm{k})$ of the flat band, similar to Fig. 1e. Note that $\Omega_{\mathrm{I}}$ does not vanish for $V\to\infty$. (c) Momentum-resolved spectral functions from DMFT at $T/t=0.025$ and $U/t=1.6$. From left to right for hybridization $V/t=0.5$, $V/t=1.5$, and $V/t=4$. (d) The imaginary part of the local self-energy at the lowest Matsubara frequency of X (red), cumulative A sites (gray), and cumulative B sites (blue) as function of $V$. The hybridization values for different group IV atoms estimated from fitting the DFT bandstructures (Fig. 3d) are indicated with arrows. “loc” and “gl” refer to the local and global energy minimum’s configuration of Ge. (e) Quantum metric in the large hybridization limit.
@@
Conclusion.  In this work, we demonstrated a topological phase transition between symmetry-distinct site-selective Mott states and a geometrically stabilized metallic state enabled by hybridization control. We find that in between two states featuring local moments in different Wyckoff positions a correlated metallic state emerges for a most simple model of one-sublattice functionalized graphene. First principles calculations demonstrate that similar physics of a site-selective Mott state transitioning to a geometrically stabilized flat-band metal can be likely realized in intercalated epitaxial graphene platforms. The series of group-IV intercalant atoms (C, Si, Ge, Sn, Pb) realize the entire range from weak ($V\ll t$) to strong hybridization ($V\gg t$).
@@
The models discussed here are indeed very similar to the topological heavy fermion description of magic angle twisted bilayer graphene [34]. In both cases, there are localized orbitals hybridizing with Dirac bands resulting in a low-energy flat band with distinct quantum geometry. Distinctly, the flat bands in chemically functionalized graphene emerge at chemical energy scales, i.e., scales of covalent bonds and corresponding hoppings tsimilar-$\sim t$. Thus, the correlated flat band physics realized here can persist potentially up to higher temperatures than in twisted bilayer graphene. It appears very promising to explore other possible emergent ordered states, like magnetism and superconductivity, and the corresponding temperature scales in these covalently functionalized graphene systems.

### v3

@@ Abstract
Figure 4: Decorated graphene with smaller impurity density. (a) Unit cell of decorated graphene with 1/8 impurity X coverage. The index of sublattice sites Ai, Bi refers to distance from the impurity X. (b) Imaginary part of the local self-energy at the lowest Matsubara frequency of X (red), cumulative A sites (gray), and cumulative B sites (blue) as function of $V$. Arrows indicate hybridization values estimated for different group IV atoms from fitting the DFT band structures (Fig. 3). Arrows for X = Ge at larger and smaller $V$ refer to the local and global energy minimum’s configuration, respectively. (c) Momentum-resolved spectral functions from DMFT at $T/t=0.025$ and $U/t=1.6$. From left to right for hybridization $V/t=0.5$, $V/t=1.5$, and $V/t=4$.
The correlated metallic state stabilized at large hybridization due to the finite quantum geometry of the flat band is similar to the vacancy-induced flat band of Ref. [35]. Building on this analogy, disorder or incomplete functionalization may redistribute spectral weight [146, 147], but the essential dichotomy between moment formation at weak $V$ and geometrically obstructed metallicity at strong $V$ is expected to persist, as long as there is no prevalent impurity dimerization [147]. Beyond our minimal model, enriching the flat band with multi-orbital degrees of freedom or spin-orbit coupling could qualitatively alter the correlated phases, for instance through Hund’s physics or topological Mott states [148, 149, 150, 151, 152, 153, 154], offering promising directions for future study. A recent theoretical proposal of transition-metal adatom systems with two flat bands and emergent Mott physics [155] offers a natural extension of such multi-orbital generalizations via graphene functionalization.
@@
Conclusion.  In this work, we demonstrated a topological phase transition between symmetry-distinct site-selective Mott states and a geometrically stabilized metallic state enabled by hybridization control. We find that in between two states featuring local moments in different Wyckoff positions a correlated metallic state emerges for a most simple model of single-sublattice functionalized graphene. First principles calculations demonstrate that similar physics of a site-selective Mott state transitioning to a geometrically stabilized flat-band metal can be likely realized in intercalated epitaxial graphene platforms. The series of group-IV intercalant atoms realize the entire range from weak ($V\ll t$) to strong hybridization ($V\gg t$).
@@
The models discussed here are indeed very similar to the topological heavy fermion description of magic angle twisted bilayer graphene [38, 111]. In both cases, localized orbitals hybridize with Dirac bands, resulting in a low-energy flat band with substantial quantum geometry. Distinctly, the flat bands in chemically functionalized graphene emerge at chemical energy scales, i.e., scales of covalent bonds and corresponding hoppings $\sim t$. Thus, the correlated flat band physics realized here can persist potentially up to higher temperatures than in twisted bilayer graphene. It appears very promising to explore other possible emergent ordered states, like magnetism and superconductivity, and the corresponding temperature scales in these covalently functionalized graphene systems.